(24 points)

  1. To be legal for certain types of prize competitions, bowling balls must be very close to 16 lb. The weights of bowling balls from a certain manufacturer are known to be normally distributed with a mean of 16 lb. and a standard deviation of .25 lb. A ball will be rejected if it weighs less than 15.68 lb.
  1. What is the probability that a ball will be rejected? Round your answer to two significant digits.

Answer:

  1. Describe how you would use a table of random digits (comprised of 0, 1, 2, 3, 4, 5, 6, 7, 8, and 9) to simulate how long it would take to identify three defective bowling balls.

Answer:

  1. Use the table below to perform the simulation you described in part B, this time for three defective bowling balls. Repeat your simulation three times, starting with the first digit of the first line and proceeding left to right, starting again at the extreme left of each successive line. What is the average waiting-time based on your three simulations?

77014

21414

95729

01392

37814

22931

94998

56569

30213

03469

16334

43057

03297

61609

68462

26199

98324

41436

96050

95744

98563

56006

93060

29402

76577

39814

75704

26127

42577

17458

25883

51840

45515

18925

46458

61380

79369

01710

93720

73046

16434

57044

23969

78022

67976

23279

67173

44918

91684

94775

Answer:

  1. What is the probability that three or more defective bowling balls are found in the first 25 examined? Do not do the calculation, but do write out the mathematical expression you would use to answer the question.

Answer:

PLEASE DON'T COPY FROM EXISTING ANSWERS, THEY ARE NOT HELPFUL
Answer #1

a)

16 = 0.25 P(x < 15.68) = P( z < (15.68-16) / 0.25) = P(z < -1.28) = 0.1003 (from Normal probability table)